# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 9 Ethical Statement, Copyright, and Fair Use Statement
Exclusion of Irrelevant Information: Do *not penalize the "Candidate Answer" for including
additional information that is not present in the "Reference Answer", *as long as that
information is medically accurate and relevant to the case*.
groundance does not affect the score: Do not penalize a response for being redundant.
@@ 4. Provide a single numerical similarity score between 0 and 100, where:
0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of
    diagnosis or supporting information. It’s completely wrong.
50: The "Candidate Answer" identifies *some relevant diagnoses or supporting information,
    but it misses major elements or includes significant inaccuracies.
75: The "Candidate Answer" correctly identifies the primary diagnosis (if applicable) and
    most of the important differential diagnoses and supporting information. It may miss some
    minor details or include some slight inaccuracies.
90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core diagnostic
    meaning and completeness. It may have minor omissions or differences in emphasis, but it
    captures almost all of the essential information.
100: The "Candidate Answer" is a perfect match in terms of core diagnostic meaning. It
    includes all of the key diagnoses and supporting information from the "Reference Answer",
    even if the wording or organization is different. Do not justify or explain output only
    the numeric score.
5. Do not justify or explain the answer. Give a score between 0 and 100, no other
text. Now, evaluate the following
Concept to Define : {Answer}
Candidate Response: {Response}
Template 3: Treatment Follow Up Prediction Prompt
As a psychiatrist, analyze the provided patient’s information, history, and follow ups to
answer the question according to the guidelines.
Guidelines
- Provide reasoning in your answer.
- Don’t use bullet points just sentences.
- No introduction before the answer just provide the answer for the provided question only.
The Patient’s History: {History}
@@
    follow = f"""First Follow Up
    {First Follow Up}
@@
    follow = f"""Second Follow up:
    {Second Follow up}
@@
Template 4: Follow Up Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text
responses about a specific mental health case, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information regarding the case assessment and
recommended follow-up contained in the second text (the "Reference Answer"). Focus on the core
clinical meaning, the completeness of the assessment understanding, and the
appropriateness/completeness of the follow-up plan, *not on writing style or specific wording.
@@
    (e.g., presenting problem, relevant history, potential diagnoses/formulations, risk
    assessment) and the core elements of the recommended follow-up plan (e.g., therapy
    recommendations, medication considerations, further assessments needed, safety planning,
    referrals).
    2. Read the "Candidate Answer". This answer will also describe an assessment of the case
    and propose a follow-up plan.
        Assessment Accuracy & Completeness: Does the "Candidate Answer" accurately reflect the
        core understanding of the patient’s situation (presenting issues, key historical factors,
        potential clinical formulation/diagnosis) as presented in the "Reference Answer"? Does
        it capture the most critical elements of the assessment?
        Follow-up Plan Completeness & Appropriateness: Does the "Candidate Answer" include the
        essential elements and intent of the follow-up plan outlined in the "Reference Answer"?
        This includes key interventions (therapy, medication), monitoring, safety considerations,
        and necessary referrals.Minor variations in terminology are acceptable if the underlying
        clinical concept and action are the same.
        Rationale/Supporting Information: Does the "Candidate Answer" implicitly or explicitly
        capture the key reasons or clinical justifications for the assessment and follow-up plan
        that are present in the "Reference Answer"? It doesn’t need every detail but should align
        on the core reasoning.
        Exclusion of Irrelevant Information: Do not penalize the "Candidate Answer" for including
        additional information *as long as that information is clinically accurate and relevant
        to the mental health case and its management*. However, significantly contradictory or
        inappropriate additions may lower the score.
        groundance does not affect the score: Do not penalize a response for being
        redundant.
@@
        0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
        of case assessment or follow-up plan. It’s completely different or clinically
        inappropriate.
        50: The "Candidate Answer" identifies some relevant aspects of the assessment or follow-
        up plan but misses major elements, contains significant inaccuracies, or proposes a
        substantially different approach.
        75: The "Candidate Answer" correctly captures the main aspects of the assessment and most
        of the essential follow-up recommendations from the "Reference Answer". It may miss some
        minor details, nuances, or justifications.
        90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core case
        understanding and the proposed follow-up plan. It may have minor omissions or differences
        in emphasis but captures almost all essential clinical information and actions.
        100: The "Candidate Answer" perfectly matches the "Reference Answer" in terms of core
        clinical meaning regarding the assessment and follow-up plan. It includes allkey
        assessment points and recommended actions, even if the wording or organization differs.
        Do not justify or explain the answer. Give a score between 0 and 100, no other text.
Now, evaluate the following
@@
Template 5: Classification Prompt
    As a psychiatrist, analyze the provided patient’s information and history to determine if
    the patient exhibits clear symptoms of any of the provided list of disorders according to
    provided guidelines.
@@
    - Concise Response: Respond with any combination of these disorders separated by a comma
    and if the poster doesn’t have any mental illness just answer with Normal.
@@
Template 6: Management Plan Prediction Prompt
You are a psychiatrist. Analyze the provided patient’s information(History, Mental State
Examination, and Physical Examination) to answer the psychiatric protocol question according
to the guidelines below.
@@
Template 7: Management Plan Evaluation Prompt
SYSTEM: You are a psychiatric protocol response evaluator. Your task is to compare two text
responses about a psychiatric protocol question, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information contained in the second text (the
"Reference Answer" or "True Label"). Focus on the clinical correctness, completeness of
@@
1. Read the "Reference Answer" carefully. Identify the key components expected in the
    psychiatric protocol response (e.g., symptom recognition, diagnostic steps, risk/safety
    considerations, treatment recommendations, ethical/legal procedures if applicable).
2. Read the "Candidate Answer". This answer will attempt to respond to the same psychiatric
    protocol question.
     Protocol Adherence & Accuracy: Does the "Candidate Answer" correctly follow the
    psychiatric protocol steps presented in the "Reference Answer"? Does it recognize key
    clinical issues, decision points, and require actions?
     Completeness of Response: Does the "Candidate Answer" address all major elements expected
    by the protocol as reflected in the "Reference Answer" (e.g., diagnostic considerations,
    intervention steps, risk management)?
     Reasoning and Justification Alignment: Does the "Candidate Answer" capture the rationale
    (explicit or implied) behind the steps taken according to the protocol in the "Reference
    Answer"? It should reflect the clinical reasoning that justifies assessment and management
    decisions.
     Handling of Additional Information: Extra information is acceptable if it is clinically
    correct and protocol-consistent. Do not penalize additional relevant detail. However,
    incorrect, contradictory, or unsafe suggestions should lower the score. groundance
    Tolerance: groundance (repeating points) does not affect the score.
@@
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of protocol steps, clinical reasoning, or appropriate action. 50: The "Candidate Answer"
    captures some relevant elements but misses major steps, has serious inaccuracies, or
    suggests significantly incorrect approaches. 75: The "Candidate Answer" captures most key
    protocol elements and clinical reasoning, with only minor omissions or inaccuracies.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of
    psychiatric protocol adherence, clinical assessment, and recommended actions, with only
    slight gaps or minor emphasis differences. 100: The "Candidate Answer" fully matches the
    "Reference Answer" in terms of psychiatric protocol steps, reasoning, and recommended
    actions, even if wording or order differ.
5. Do not justify or explain the answer. Only provide a single numerical score between 0 and
100. No other text output.Now, evaluate the following
@@
Template 8: Clinical approach Prompt
Patient History: {History}
@@
Generated Question: Generate exactly one clear, focused, and medically relevant question
based on Clinical Approach.
@@
You are an expert grader evaluating whether a history question is relevant to a given
clinical approach.
@@
Determine if the question contains keywords or semantic meaning that are directly related to
the clinical approach. Respond only with "yes" if it is relevant or "no" if it is not. Do
not provide any explanation.
@@
You are a psychiatrist. Based on the patient’s clinical history and the
question, provide a clear, accurate, and medically appropriate answer.
Patient History:{History}
@@
You are a clinical reasoning expert. Based on the following patient case history, write a
concise summary of the clinical approach. Include possible diagnoses, recommended
investigations, and management steps. The tone should be similar to a medical case
explanation, like those found in exam review books.
Patient History:{History}
Write a clinical approach (1-2 paragraphs):
Clinical Approach Prediction Prompt
Template 9: Clinical Approach Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text
responses about a specific mental health case, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information regarding the clinical approach
contained in the second text (the "Reference Answer"). Focus on the clinical reasoning,
assessment strategy, and diagnostic insight not on writing style or specific wording.
@@
1. Read the "Reference Answer" carefully. Identify the key components of the clinical
    approach: presenting symptoms, history taking, differential diagnosis, initial
    evaluation steps, and rationale for assessment strategy.
2. Read the "Candidate Answer". This is another attempt to outline a clinical approach to
    the same case.
@@ 3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
    Assessment Strategy Accuracy & Completeness: Does the "Candidate Answer" reflect an
    accurate and complete approach toassessing the case?
    Does it capture the key symptoms, relevant history, and potential clinical hypotheses
    found in the "Reference Answer"?
    Diagnostic Reasoning & Thought Process: Does the "Candidate Answer" follow a similar
    line of clinical reasoning, including formulation or differential diagnosis as in the
    "Reference Answer"?
    Clinical Actions or Evaluation Steps: Does it propose similar or equivalent steps in the
    evaluation process (e.g., screeningtools, physical exams, lab tests, psychiatric
    interview elements)?
@@
    justifications provided in the "Reference Answer"? Perfect wording is not needed focus
    on reasoning.
    No Penalty for Relevance: Do not penalize clinically valid additions if they are
    appropriate, even if not found in the"Reference Answer". groundance does not grace score.
@@ 4. Provide a single numeric score between 0 and 100:
Do not justify or explain output only the numeric score.
Now, evaluate the following
@@
Template 10: Three-stage Mental QA prompt pipeline. (a) Prompt 1 directs the model to answer clinical questions step-by-step using patient history in a structured, professional tone; (b) Prompt 2 refocuses on defining the target mental-health concept clearly and concisely, as in exam-style review; (c) Evaluation Prompt instructs an automated scorer to assign a 0-100 accuracy score based on completeness, clarity, depth, and relevance of the model’s definition.
You are a psychiatrist specializing in mental health knowledge. Based on the provided dataset,
answer the following question accurately
and concisely. Ensure your response is well-structured, using bullet points or numbered lists
where appropriate, and maintain a professional
tone throughout. Think step by step to provide a logical and coherent answer like those found
in exam review books.
History: {History}
@@
Mental QA Prediction Prompt 1
You are an expert psychiatrist with deep knowledge of mental health concepts. Your task is to
use the patient’s history to answer the
clinical question by defining or explaining the target concept clearly and concisely like
those found in exam review books.
History: {History}
Question: {Question}
Mental QA Prediction Prompt 2
System: You are a mental health concept knowledge evaluator. Your task is to assess how
accurately, completely, and clearly the candi-
date’s response defines the concept provided in the "Answer" field, taking into account the
clinical context in the "History."
@@
    1. Read the "Answer" this is the clinical concept or term to define (e.g., "Loss of
    interest or pleasure in activities. . . ").

### v2

@@ 3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
     Information: Do *not penalize the "Candidate Answer" for including additional, medically accurate and
     relevant information not found in the "Reference Answer".groundance does not affect the score:
     Do not penalize a response for being redundant.
@@ 4. Provide a single numerical similarity score between 0 and 100, where:
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of treatment
     or rationale. It is completely inappropriate or unrelated.
     50: The "Candidate Answer" identifies *some relevant treatment components or strategies, but misses
     key parts or includes significant inaccuracies.
     75: The "Candidate Answer" correctly identifies the primary treatment and most important additional
     strategies or reasoning. Minor omissions or slight inaccuracies are acceptable.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core treatment strategies
     and rationale. It may have minor differences in emphasis or detail.
     100: The "Candidate Answer" is a perfect match in terms of treatment content and rationale.
     It includes all key interventions and supporting reasoning from the "Reference Answer",
     even if the wording or structure is different.
5.Do not justify or explain the answer.Give a score between 0 and 100, no other text.Now evaluate the following
Reference Answer:{Answer}
Candidate Answer:{Response}
Template 5: Treatment Follow Up Prediction Prompt
As a psychiatrist, analyze the provided patient’s information, history, and follow ups to answer the question according to the guidelines. Guidelines - Provide reasoning in your answer. - Don’t use bullet points just sentences. - No introduction before the answer just provide the answer for the provided question only. The Patient’s History: {History}
@@
    follow = f"""First Follow Up {First Follow Up}"""
@@
    follow = f"""Second Follow up:{Second Follow up}"""
@@
Template 6: Treatment Follow Up Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text responses about a
specific mental health case, assessing the degree to which the first text (the "Candidate Answer") captures
the essential information regarding the case assessment and recommended follow-up contained in the second text
(the "Reference Answer"). Focus on the core clinical meaning, the completeness of the assessment understanding,
and the appropriateness/completeness of the follow-up plan, *not on writing style or specific wording.
@@
    (e.g., presenting problem, relevant history, potential diagnoses/formulations, risk assessment) and the
    core elements of the recommended follow-up plan (e.g., therapy recommendations, medication considerations,
    further assessments needed, safety planning, referrals).
    2. Read the "Candidate Answer". This answer will also describe an assessment of the case and propose a
    follow-up plan.
    3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:Assessment
    Accuracy & Completeness: Does the "Candidate Answer" accurately reflect the core understanding of the
    patient’s situation (presenting issues, key historical factors, potential clinical formulation/diagnosis)
    as presented in the "Reference Answer"? Does it capture the most critical elements of the assessment?
    Follow-up Plan Completeness & Appropriateness: Does the "Candidate Answer" include the essential elements
    and intent of the follow-up plan outlined in the "Reference Answer"? This includes key interventions
    (therapy, medication), monitoring, safety considerations, and necessary referrals.Minor variations in
    terminology are acceptable if the underlying clinical concept and action are the same. Rationale/Supporting
    Information: Does the "Candidate Answer" implicitly or explicitly capture the key reasons or clinical
    justifications for the assessment and follow-up plan that are present in the "Reference Answer"? It doesn’t
    need every detail but should align on the core reasoning.Exclusion of Irrelevant Information:
    Do not penalize the "Candidate Answer" for including additional information *as long as that information is
    clinically accurate and relevant to the mental health case and its management*. However, significantly
    contradictory or inappropriate additions may lower the score.groundance does not affect the score:
    Do not penalize a response for being redundant.
@@
        0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of case
        assessment or follow-up plan. It’s completely different or clinically inappropriate.
        50: The "Candidate Answer" identifies some relevant aspects of the assessment or follow-up plan but misses
        major elements, contains significant inaccuracies, or proposes a substantially different approach.
        75: The "Candidate Answer" correctly captures the main aspects of the assessment and most of the essential
        follow-up recommendations from the "Reference Answer". It may miss some minor details, nuances, or
        justifications.
        90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core case understanding and
        the proposed follow-up plan. It may have minor omissions or differences in emphasis but captures almost
        all essential clinical information and actions.
        100: The "Candidate Answer" perfectly matches the "Reference Answer" in terms of core clinical meaning
        regarding the assessment and follow-up plan. It includes all key assessment points and recommended actions
        , even if the wording or organization differs. Do not justify or explain the answer. Give a score between
        0 and 100, no other text.Now, evaluate the following
@@
Template 7: Classification Prompt
    As a psychiatrist, analyze the provided patient’s information and history to determine if the patient
    exhibits clear symptoms of any of the provided list of disorders according to provided guidelines.
@@
    - Concise Response: Respond with any combination of these disorders separated by a comma and if the poster
    doesn’t have any mental illness just answer with Normal.
@@
Template 8: Management Plan Prediction Prompt
You are a psychiatrist. Analyze the provided patient’s information(History, Mental State Examination,
and Physical Examination) to answer the psychiatric protocol question according to the guidelines below.
@@
Template 9: Management Plan Evaluation Prompt
SYSTEM: You are a psychiatric protocol response evaluator. Your task is to compare two text responses about a
psychiatric protocol question, assessing the degree to which the first text (the "Candidate Answer") captures
the essential information contained in the second text (the "Reference Answer" or "True Label"). Focus on
@@
1. Read the "Reference Answer" carefully. Identify the key components expected in the psychiatric protocol
response (e.g., symptom recognition, diagnostic steps, risk/safety considerations, treatment recommendations
ethical/legal procedures if applicable).
2. Read the "Candidate Answer". This answer will attempt to respond to the same psychiatric protocol question.
3. Compare the "Candidate Answer" to the "Reference Answer" based on the
following criteria:
     Protocol Adherence & Accuracy: Does the "Candidate Answer" correctly follow the psychiatric protocol steps
     presented in the "Reference Answer"? Does it recognize key clinical issues, decision points, and require
     actions?  Completeness of Response: Does the "Candidate Answer" address all major elements expected by the
     protocol as reflected in the "Reference Answer" (e.g., diagnostic considerations, intervention steps,
     risk management)? Reasoning and Justification Alignment: Does the "Candidate Answer" capture the rationale
     (explicit or implied) behind the steps taken according to the protocol in the  "Reference Answer"? It should
     reflect the clinical reasoning that justifies  assessment and management  decisions.
     Handling of Additional Information: Extra information is acceptable if it is clinically correct and
     protocol-consistent. Do not penalize additional relevant detail. However, incorrect, contradictory, or
     unsafe suggestions should lower the score. groundance Tolerance: groundance (repeating points) does
     not affect the score.
@@
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of protocol steps,
     clinical reasoning, or appropriate action.
     50: The "Candidate Answer" captures some relevant elements but misses major steps, has serious inaccuracies
     or suggests significantly incorrect approaches.
     75: The "Candidate Answer" captures most key protocol elements and clinical reasoning, with only minor
     omissions or inaccuracies.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of psychiatric protocol
     adherence, clinical assessment, and recommended actions,  with only slight gaps or minor emphasis
     differences.
     100: The "Candidate Answer" fully matches the "Reference Answer" in terms of psychiatric protocol steps,
     reasoning, and recommended actions, even if wording or order differ.
5. Do not justify or explain the answer. Only provide a single numerical score between 0 and 100.
No other text output.Now, evaluate the following
@@
Template 10: Clinical approach Prompt
The Patient’s History: {History}
@@
Generated Question: Generate exactly one clear, focused, and medically relevant question based on
Clinical Approach.
@@
You are an expert grader evaluating whether a history question is relevant to a given clinical approach.
@@
Determine if the question contains keywords or semantic meaning that are directly related to the clinical
approach. Respond only with "yes" if it is relevant or "no" if it is not. Do not provide any explanation.
@@
You are a psychiatrist. Based on the patient’s clinical history and the question, provide a clear, accurate,
and medically appropriate answer.
The Patient’s History:{History}
@@
Template 11: Clinical Approach Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text responses about a
specific mental health case, assessing the degree to which the first text (the "Candidate Answer") captures
The essential information regarding the clinical approach is contained in the second text("Reference Answer")
Focus on the clinical reasoning, assessment strategy, and diagnostic insight, not on writing style .
@@
1. Read the "Reference Answer" carefully. Identify the key components of the clinical approach: presenting
symptoms, history taking, differential diagnosis, initial evaluation steps, and rationale for
assessment strategy.
2. Read the "Candidate Answer". This is another attempt to outline a clinical approach to the same case.
@@ 3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
    Assessment Strategy Accuracy & Completeness: Does the "Candidate Answer" reflect an accurate and complete
    approach toassessing the case? Does it capture the key symptoms, relevant history, and potential clinical
    hypotheses found in the "Reference Answer"? Diagnostic Reasoning & Thought Process: Does the
    "Candidate Answer" follow a similar line of clinical reasoning, including formulation or differential
    diagnosis as in the "Reference Answer"? Evaluation Steps: Does it propose similar or equivalent steps in
    the evaluation process (e.g., screeningtools, physical exams, lab tests, psychiatric interview elements)?
@@
    justifications provided in the "Reference Answer"? Perfect wording is not needed focus on reasoning.No
    Penalty for Relevance: Do not penalize clinically valid additions if they are appropriate, even if not found
    in the"Reference Answer". groundance does not grace score.
@@ 4. Provide a single numeric score between 0 and 100:
    Do not justify or explain output only the numeric score. Now, evaluate the following
@@
Template 12: Two-stage Mental QA prompt pipeline. (a) Prediction Prompt directs the model to define the target mental-health concept clearly and concisely, as in an exam-style review, emphasizing conceptual precision and professional tone; (b) Evaluation Prompt instructs an automated scorer to assign a 0–100 accuracy score based on the completeness, clarity, depth, and relevance of the model’s definition.
You are a psychiatrist specializing in mental health knowledge. Based on the provided dataset, answer the
following question accurately and concisely. Ensure your response is well-structured, using bullet points
or number lists where appropriate, and maintain a professional tone throughout. Think step by step to provide
a logical and coherent answer like those found in exam review books.
The Patient’s History: {History}
@@
Mental QA Prediction Prompt
System: You are a mental health concept knowledge evaluator. Your task is to assess how accurately,completely,
and clearly, the candidate’s response defines the concept provided in the "Answer" field, taking into account
the clinical context in the "History."
@@
    1. Read the "Answer" this is the clinical concept or term to define
    (e.g., "Loss of interest or pleasure in activities").